1.Differentiate between real and non-real numbers (give an examples).
2. Differentiate between rational and irrational numbers (give examples).
1
Expert's answer
2020-05-07T16:05:54-0400

1. Real numbers are all numbers that we can be displayed in a number line.

example

"-3, -1, 0, 1, \sqrt{2}, \pi, e."

All others are non-real numbers. They are called "imaginary" numbers.

example

"\sqrt{-2}, 4-3i"

2. Rational numbers are fractions.

Example:

"\frac{3}{5}, -\frac{6}{8}, 5\frac{2}{9}"

Irrational numbers are numbers that cannot be written as a rational fraction because they are infinite.

"\sqrt2, \pi, \sqrt3."